# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ IV-B Geometric construction of the flow and jump sets
When the robot operates in the move-to-target mode, its velocity is directed towards the target location. Hence, if the path joining the robot’s location and the target is obstructed by an obstacle, let’s say $\mathcal{O}_{i}$ for some $i\in\mathbb{I}$ i.e., $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$, then the center of the robot enters the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{i}$ through the landing region $\mathcal{R}_{l}^{i},$ which is defined as
$\mathcal{R}_{l}^{i}:=\left\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))|{\mathcal{L}_{s}(\mathbf{x},\mathbf{0})}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset\right\}.$ (15)
The union of the landing regions over all obstacles is defined as
$\mathcal{R}_{l}:=\bigcup_{i\in\mathbb{I}}\mathcal{R}_{l}^{i}.$ (16)
@@
Next, we define an exit region $\mathcal{R}_{e}$ as the part of the $\gamma-$neighborhood of the $r_{a}-$dilated obstacles that do not belong the landing region, as shown in Fig. 1b. The exit region $\mathcal{R}_{e}$ is defined as
$\mathcal{R}_{e}=\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus\mathcal{R}_{l}.$ (17)
@@
Note that when the robot’s center at $\mathbf{x}$ belongs to the exit region, the line segment connecting its location to the target location $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})$ does not intersect with the interior of the nearest $r_{a}-$dilated obstacle. Hence, the robot should move straight towards the target location only if it is in the exit region. Next, we provide the geometric construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (11).
@@ IV-B1 Flow and jump sets (move-to-target mode)
When the robot operating in the move-to-target mode, enters in the landing region $\mathcal{R}_{l}$ (16), it should switch to the obstacle-avoidance mode to avoid collision. Hence, the jump set of the move-to-target mode for the state $\mathbf{x}$ is defined as
$\mathcal{J}_{0}^{\mathcal{W}}:=\overline{\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{R}_{l}},$ (18)
where $\gamma_{s}\in(0,\gamma).$ For robustness purposes (with respect to noise), we introduce a hysteresis region by allowing the robot, operating in the move-to-target mode inside the $(r_{a}+\gamma)-$neighborhood of the set $\mathcal{O}_{\mathcal{W}}$, to move closer to the set $\mathcal{O}_{\mathcal{W}}$ before switching to the obstacle-avoidance mode.
@@
$\mathcal{F}_{0}^{\mathcal{W}}:=\left(\mathcal{W}\setminus(\mathcal{D}_{r_{a}+\gamma_{s}}^{\circ}(\mathcal{O}_{\mathcal{W}}))\right)\cup\mathcal{R}_{e}.$ (19)
Notice that the union of the jump set $\mathcal{J}_{0}^{\mathcal{W}}$ and the flow set $\mathcal{F}_{0}^{\mathcal{W}}$ covers the robot-centered obstacle-free workspace $\mathcal{W}_{r_{a}}$.
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
The robot operates in the obstacle-avoidance mode inside the $\gamma-$neighborhood of the obstacle-occupied workspace. Since the robot can safely move straight towards the target location and exit the $\gamma-$neighborhood of the nearest obstacle, whenever it is in the exit region (17), it should switch back to the move-to-target mode only if it is in the exit region. To that end, we make use of the hit point $\mathbf{h}$ (i.e., the location of the center of the robot when it switched from the move-to-target mode to the current obstacle-avoidance mode) to define the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}^{\mathcal{W}}$ for the state $\mathbf{x}$ as follows:
$\mathcal{J}_{1}^{\mathcal{W}}:=\left(\mathcal{W}\setminus\mathcal{D}_{r_{a}+\gamma}^{\circ}(\mathcal{O}_{\mathcal{W}})\right)\cup\mathcal{ER}^{\mathbf{h}}{\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))}\cup\mathcal{S}_{\mathbf{0}},$ (21)
where the set $\mathcal{S}_{\mathbf{0}}:=\overline{\mathcal{B}_{\gamma}(\mathbf{0})\setminus\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})}$. The inclusion of the set $\mathcal{S}_{\mathbf{0}}$ in the set $\mathcal{J}_{1}^{\mathcal{W}}$ allows us to ensure the stability of the origin, as stated later in Theorem 1.
For a hit point $\mathbf{h}\in\mathcal{W}_{r_{a}}$, the set $\mathcal{ER}^{\mathbf{h}}$ is given by
@@
with $\epsilon\in(0,\bar{\epsilon}]$ where $\bar{\epsilon}$ is a sufficiently small positive scalar. The set $\mathcal{ER}^{\mathbf{h}}$ contains the locations from the exit region $\mathcal{R}_{e}$ which are at least $\epsilon$ units closer to the target location than the current hit point $\mathbf{h}$. Since the obstacles are compact and convex, and the target location $\mathbf{0}$ is within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}$, it is possible to guarantee the existence of the parameter $\bar{\epsilon}$ such that the intersection set $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ is non-empty for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for each $i\in\mathbb{I},$ as stated in the following lemma:
@@
Let Assumption 8 hold. Then, for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, there exists $\bar{\epsilon}>0$ such that for any $\epsilon\in(0,\bar{\epsilon}]$ the set $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset,$ where the set $\mathcal{ER}^{\mathbf{h}}$ is defined in (22).
@@
See Appendix IX-A. ∎
According to (21) and (22), the robot operating in the obstacle-avoidance mode, can switch to the move-to-target mode when its center belongs to the exit region $\mathcal{R}_{e}$ and the target location $\mathbf{0}$ is at least $\epsilon$ units closer to $\mathbf{x}$ than the current hit point $\mathbf{h}$.
The flow set of the obstacle-avoidance mode for the state $\mathbf{x}$ i.e., $\mathcal{F}_{1}^{\mathcal{W}}$ is defined as follows:
$\mathcal{F}_{1}^{\mathcal{W}}:=\overline{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus{(\mathcal{ER}^{\mathbf{h}}\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0})))}}.$ (23)
Notice that the union of the jump set (21) and the flow set (23) exactly covers the robot-centered obstacle-fee workspace $\mathcal{W}_{r_{a}}$.
@@
Given that the workspace $\mathcal{W}$ is both convex and compact, there may exist some locations $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from which the nearest point to the robot’s center on the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ is not unique. This scenario prevents the implementation of the obstacle-avoidance term $\mathbf{v}(\mathbf{x},\mathbf{a})$ in the control law, at such locations. However, by excluding the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from the set $\mathcal{F}_{1}^{\mathcal{W}}$, as defined in (23), it is ensured that the obstacle-avoidance mode is never activated within the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$.
@@
where $\mathcal{J}_{\mathbf{x}}:=\{\xi\in\mathcal{K}|m=1,\mathbf{x}\in\mathcal{J}_{1}^{\mathcal{W}}\}$ and $\mathcal{J}_{s}:=\{\xi\in\mathcal{K}|m=1,s=[s_{0},s_{0}+\delta_{s}]\}$ with $s_{0}\in\mathbb{R}_{\geq 0}$ being the initial value of the state $s$, i.e., $s(0,0)=s_{0}$, and $0<\delta_{s}<\tau_{s}$ for some $\tau_{s}>0$.
@@
The definition of sets in (24) enables the control to immediately switch to the move-to-target mode if it is initialized in the obstacle-avoidance mode (i.e., $\xi(0,0)\in\mathcal{J}_{1}$). This ensures that the hit point $\mathbf{h}$ always belongs to the set $\mathcal{J}_{0}^{\mathcal{W}}$ before the robot starts moving in the obstacle-avoidance mode, thus guaranteeing the existence of the parameter $\bar{\epsilon}$, as stated in Lemma 1.

### v5

@@ IV-B Geometric construction of the flow and jump sets
When the robot is located at a distance larger than $\gamma$ from the obstacles, it moves straight towards the target location in the move-to-target mode. The robot’s distance from the nearby obstacles is obtained by evaluating $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})$, as discussed in Section II-B, where the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ is defined in (8). Upon entering in the $\gamma$-neighbourhood of the obstacles, as per Assumption 9, the robot is within the $\gamma$-neighbourhood of only one obstacle. In other words, according to Assumption 9, the fact that $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma$, implies the existence of $i\in\mathbb{I}$ such that $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and $\mathbf{x}\notin\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{j}))$ for all $j\in\mathbb{I}\setminus{i}$.
Figure 3: A workspace scenario in which condition (18) is satisfied.
Figure 4: Partitioning of the region $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, with $r_{a}=0.25\,m$ and $\gamma=1.15\,m$, into the avoidance region $\mathcal{R}_{a}^{i}$ and the exit region $\mathcal{R}_{e}^{i}$, for different values of the sensing radius $R_{s}$: (a) $R_{s}=0.5\,m$, (b) $R_{s}=1\,m$, (c) $R_{s}=1.5\,m$, (d) $R_{s}=4\,m$.
@@
When the robot operates in the move-to-target mode, its velocity is directed towards the target location. Hence, if the robot enters in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$, it should constantly verify whether the path joining the robot’s location and the target is obstructed by $\mathcal{O}_{i}$. Therefore, we construct the set ð$\eth_{i}^{\mathbf{x}}$, which contains locations on the boundary of obstacle $\mathcal{O}_{i}$ that have a clear line of sight to the center of the robot. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, the set ð$\eth_{i}^{\mathbf{x}}$ is defined as
ð$\eth_{i}^{\mathbf{x}}:=\{\mathbf{y}\in\eth_{\mathcal{W}}^{\mathbf{x}}\,|\,\mathbf{y}\in\partial\mathcal{O}_{i}\}.$ (17)
Observe that if $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ and
ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset,$ (18)
as shown in Fig. 3, then the robot, moving straight towards the target location, will eventually collide with obstacle $\mathcal{O}_{i}$, where $\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$ represents the cylinder with width $2r_{a}$ and axis $\mathcal{L}_{s}(\mathbf{x},\mathbf{0})$ as defined in Section II-C5. We define the avoidance region, denoted by $\mathcal{R}_{a}^{i}$, as the set of all locations from the $\gamma-$neighbourhood of $r_{a}-$dilated obstacle $\mathcal{O}_{i}$ such that condition (18) is satisfied. The set $\mathcal{R}_{a}^{i}$ is defined as
$\mathcal{R}_{a}^{i}:=\left\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))|\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset\right\}.$ (19)
Since the set ð$\eth_{i}^{\mathbf{x}}$ is used to define the avoidance region $\mathcal{R}_{a}^{i}$, the shape of $\mathcal{R}_{a}^{i}$ depends on the value of the sensing radius $R_{s}$, as shown in Fig. 4. However, as stated in the next lemma, irrespective of the value of the sensing radius $R_{s}$, the avoidance region $\mathcal{R}_{a}^{i}$ is always a subset of the landing region $\mathcal{R}_{l}^{i}$, which is defined as
$\mathcal{R}_{l}^{i}:=\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))|\mathcal{L}_{s}(\mathbf{x},\mathbf{0})\cap$ $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\neq\emptyset,$ (20)
$\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0\}.$
Furthermore, if $R_{s}>l_{i}$, then one has $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$, where $l_{i}$ is the largest possible distance between any two points from the set $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, that is
$l_{i}=\max\{\lVert\mathbf{p}-\mathbf{q}\rVert|\mathbf{p},\mathbf{q}\in\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})\}.$ (21)
Lemma 1.
For each $i\in\mathbb{I}\setminus\{0\}$ and for any $R_{s}>0$, it holds that $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$. Additionally, if $R_{s}>l_{i}$, where $l_{i}$ is defined in (21), then $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$.
Proof.
See Appendix -A. ∎
In Fig. 4, the blue curve represents the boundary of the landing region $\mathcal{R}_{l}^{i}$ for obstacle $\mathcal{O}_{i}$. Notice that the avoidance region $\mathcal{R}_{a}^{i}$ is always a subset of the landing region $\mathcal{R}_{l}^{i}$, irrespective of the value of the sensing radius $R_{s}$.
Since $\mathcal{W}$ is a closed convex set, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$, condition (18) is not satisfied, where $\mathcal{O}_{0}=\left(\mathcal{W}^{\circ}\right)^{c}$. Therefore, $\mathcal{R}_{a}^{0}$ is an empty set. The union of the avoidance regions over all obstacles is given by
$\mathcal{R}_{a}:=\bigcup_{i\in\mathbb{I}}\mathcal{R}_{a}^{i}.$ (22)
@@
Next, we define the exit region $\mathcal{R}_{e}$ as the part of the $\gamma-$neighbourhood of the $r_{a}-$dilated obstacles that do not belong the avoidance region. The exit region $\mathcal{R}_{e}$ is defined as
$\mathcal{R}_{e}=\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus\mathcal{R}_{a}.$ (23)
@@
Note that when the robot’s center $\mathbf{x}$ belongs to the exit region, condition (18) is not satisfied. In other words, if the robot, with its center in the exit region, moves directly towards the target location along the path connecting its center to the origin, it will not collide with the nearest obstacle within the sensing radius $R_{s}$. Hence, the robot should move straight towards the target location only if it is in the exit region.
Next, we provide the geometric construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, used in (12).
@@ IV-B1 Flow and jump sets (move-to-target mode)
When the robot, in the move-to-target mode, enters the $\gamma$-neighborhood of an obstacle obstructing its straight path to the target location, the control input switches to the obstacle-avoidance mode. Hence, the jump set of the move-to-target mode for the state $\mathbf{x}$ is defined as
$\mathcal{J}_{0}^{\mathcal{W}}:={\mathcal{N}_{\gamma_{s}}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{R}_{a}},$ (24)
where $\gamma_{s}\in(0,\gamma).$ In (24), we allow the robot to enter the $\gamma_{s}$-neighborhood of obstacles before switching to the obstacle-avoidance mode. This creates a hysteresis region, $\mathcal{N}_{\gamma-\gamma_{s}}(\mathcal{D}_{r_{a}+\gamma_{s}}(\mathcal{O}_{\mathcal{W}}))$, which acts as a buffer zone to prevent frequent switching between the modes due to small disturbances or noise.
@@
$\mathcal{F}_{0}^{\mathcal{W}}:=\left(\mathcal{W}\setminus(\mathcal{D}_{r_{a}+\gamma_{s}}^{\circ}(\mathcal{O}_{\mathcal{W}}))\right)\cup\overline{\mathcal{R}_{e}}.$ (25)
Notice that the union of the jump set $\mathcal{J}_{0}^{\mathcal{W}}$ and the flow set $\mathcal{F}_{0}^{\mathcal{W}}$ covers the obstacle-free workspace $\mathcal{W}_{r_{a}}$, as shown in Fig. 5.
@@
Figure 5: Two-dimensional illustration of the flow sets and the jump sets considered in Sections IV-B1 and IV-B2.
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
In the obstacle-avoidance mode, the robot moves along the boundary of the obstacle being avoided in its $\gamma$-neighborhood. The control can switch back to the move-to-target mode only if the obstacle no longer obstructs the robot’s straight path to the origin. Thus, the switching occurs only when $\mathbf{x}$ is in the exit region $\mathcal{R}_{e}$ (23). Additionally, the hit point $\mathbf{h}$ (the robot’s position when it switched from the move-to-target to the current obstacle-avoidance mode) is used to ensure that, when the control input switches back to the move-to-target mode, the target location at the origin is closer to $\mathbf{x}$ than to $\mathbf{h}$. This condition helps us to establish the global convergence properties of the target location, as stated in Theorem 1. The jump set for the obstacle-avoidance mode $\mathcal{J}_{1}^{\mathcal{W}}$ is defined as follows:
$\mathcal{J}_{1}^{\mathcal{W}}:=\left(\mathcal{W}\setminus\mathcal{D}_{r_{a}+\gamma}^{\circ}(\mathcal{O}_{\mathcal{W}})\right)\cup\overline{\mathcal{ER}^{\mathbf{h}}}\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))\cup\mathcal{S}_{\mathbf{0}}.$ (27)
For any hit point $\mathbf{h}\in\mathcal{W}_{r_{a}}$, the set $\mathcal{ER}^{\mathbf{h}}$ is given by
@@
with $\epsilon\in(0,\bar{\epsilon}]$, where $\bar{\epsilon}$ is a sufficiently small positive scalar. The set $\mathcal{ER}^{\mathbf{h}}$ contains the locations $\mathbf{x}$ from the exit region $\mathcal{R}_{e}$ for which the target location is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$. Since the obstacles are compact and convex, and the target location $\mathbf{0}$ is within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}$, it is possible to guarantee the existence of the parameter $\bar{\epsilon}$ such that the intersection set $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ is non-empty for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for each $i\in\mathbb{I}\setminus\{0\}$, as stated in the following lemma:
@@
Let Assumption 9 hold. Then, for each $i\in\mathbb{I}\setminus\{0\}$, for every $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, there exists $\bar{\epsilon}>0$ such that for any $\epsilon\in(0,\bar{\epsilon}]$, one has $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset,$ where $\mathcal{ER}^{\mathbf{h}}$ is defined in (28).
@@
See Appendix -B. ∎
According to (27) and (28), the control input corresponding to the obstacle-avoidance mode, can switch to the move-to-target mode when $\mathbf{x}$ belongs to the exit region $\mathcal{R}_{e}$ and the target location $\mathbf{0}$ is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$. This introduces a hysteresis region that prevents frequent switching between the modes due to small disturbances or noise.
Next, we define the set $\mathcal{S}_{0}$, used in (27).
Figure 6: (a) Illustration of the region $\mathcal{J}1^{\mathcal{W}}\setminus\mathcal{S}_{\mathbf{0}}$. (b) Illustration of the region $\mathcal{J}_{1}^{\mathcal{W}}$.
The set $\mathcal{S}_{\mathbf{0}}$, which contains locations from the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{B}_{R_{s}}(\mathbf{0})$ from which the robot can safely move straight towards the target location at the origin, is given by
$\mathcal{S}_{\mathbf{0}}:=\{\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\cap\mathcal{B}_{R_{s}}(\mathbf{0})|\eth_{\mathcal{W}}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset\}.$ (29)
Notice that if $\mathcal{S}_{\mathbf{0}}$ is excluded from the set $\mathcal{J}_{1}^{\mathcal{W}}$ in (27) and $\mathbf{0}$ is in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$ for some $i\in\mathbb{I}\setminus\{0\}$, then the $\mathbf{x}$ trajectory, starting from some $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, may not enter in the set $\overline{\mathcal{ER}^{\mathbf{h}}}$ and as such in the set $\mathcal{J}_{1}^{\mathcal{W}}$ when $\epsilon$ is set relatively high. This may cause the robot to indefinitely operate in the obstacle-avoidance mode in the $\gamma-$neighbourhood of obstacle $\mathcal{O}_{i}$. For example, consider Fig. 6a, in which the $\mathbf{x}$ trajectory starting from $\mathbf{h}$ does not enter the set $\mathcal{J}_{1}^{\mathcal{W}}\setminus\mathcal{S}_{\mathbf{0}}$. Therefore, we include the set $\mathcal{S}_{\mathbf{0}}$ in (27) to ensure that the $\mathbf{x}$ trajectories starting from any $\mathbf{h}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ always enter in the set $\mathcal{J}_{1}^{\mathcal{W}}$ when the control input corresponds to the obstacle-avoidance mode, as shown in Fig. 6b.
The flow set of the obstacle-avoidance mode for the state $\mathbf{x}$, i.e., $\mathcal{F}_{1}^{\mathcal{W}}$, is defined as follows:
$\mathcal{F}_{1}^{\mathcal{W}}:=\overline{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))\setminus{(\mathcal{ER}^{\mathbf{h}}\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))\cup\mathcal{S}_{\mathbf{0}})}}.$ (30)
Notice that the union of the jump set (27) and the flow set (30) exactly covers the obstacle-free workspace $\mathcal{W}_{r_{a}}$, as shown in Fig. 5.
@@
Given that the workspace $\mathcal{W}$ is both convex and compact, there may exist some locations $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from which the nearest point to the robot’s center on the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ is not unique. This scenario prevents the implementation of the obstacle-avoidance term $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in the control law, at such locations. However, by excluding the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$ from the set $\mathcal{F}_{1}^{\mathcal{W}}$, as defined in (30), it is ensured that the obstacle-avoidance mode is never activated within the set $\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))$.
@@
where $\mathcal{J}_{\mathbf{x}}:=\{\xi\in\mathcal{K}|m=1,\mathbf{x}\in\mathcal{J}_{1}^{\mathcal{W}}\}$ and $\mathcal{J}_{s}:=\{\xi\in\mathcal{K}|m=1,s\in[s_{0},s_{0}+\delta_{s}]\}$ with $s_{0}\in\mathbb{R}_{\geq 0}$ being the initial value of the state $s$, i.e., $s(0,0)=s_{0}$, and $0<\delta_{s}<\tau_{s}$ for some $\tau_{s}>0$.
@@
The definition of sets in (31) enables the control to immediately switch to the move-to-target mode if it is initialized in the obstacle-avoidance mode (i.e., $\xi(0,0)\in\mathcal{J}_{1}$). This ensures that the hit point $\mathbf{h}$ always belongs to the set $\mathcal{J}_{0}^{\mathcal{W}}$ before the robot starts moving in the obstacle-avoidance mode, thereby confirming the existence of the parameter $\bar{\epsilon}$, as established in Lemma 2.